which line is perpendicular to the line whose equation is 5y+6=-3x ?
Answer by Alan3354(69443)

which line is perpendicular to the line whose equation is 5y+6=-3x ?
y = (-3/5)x - 6/5
Slope = -3/5
-------------
The slope of lines perpendicular is the negative inverse = 5/3
------
y = (5/3)x + b, b is any real number